Question 5. Suppose we have 3 point charges fixed in place in a line as below:



P ............Q.................-Q................-2Q. ( assume equidistance between)

What is the total electric potential due to these three charges at point P?

a.
keQ/6r

b.
-2keQ/r

c.
keQ/3r
d.
0O

e.
-2keQ/3r
Answer #1


Electric potential is V = k[(Q/r)+(-Q/2r)-(2Q/3r)]

V = (k*Q/r)(-1/6) = -k*Q/6r


so option a is correct
Answer #2

Electric Potential is given as

V = k Q/r

Q = charge   , r = distance from charge

Potential by charge Q at distance ''r'' from point P:

V1 = k Q /r

Potential by charge -Q at distance ''2r'' from point P:

V2 = -k Q /2r

Potential by charge -2Q at distance ''3r'' from point P:

V3 = -k 2Q /3r

Total Electric Potential :

Vtotal = V1 + V2 + V3\

Vtotal = (k Q /r )+ (-k Q /2r) + ( -k 2Q /3r)

Vtotal = -kQ/6r